Tdickinson71:

A salt truck drops 39 kilograms of salt per minut. How many grams of salt does the truck drop per second

Nnesha:

conversation factor 1 minute =60 second `39 kilograms of salt per minute` => \[\rm 39\frac{kg}{minute}\]\[\rm 39\frac{kg} {minute} \cdot \frac{1~ minute}{60 ~second}\] we are trying to get rid of minutes, therefore the minutes will be at the numerator.